# Linear constellations in primes with arithmetic restrictions

### v2

@@ 11.3. Three-term arithmetic progressions in primes with prescribed primitive roots
Apart from the application of the new Proposition A.2, this is essentially the same as the proof of [24, Theorem 5.5], so we will be brief and focus on what is different.
We consider the splitting field $E/\mathbb{Q}(t)$ of the polynomial $f(t,X)=X^{n}-t(X-1)$, which is ramified over the places $t\mapsto 0,t\mapsto\infty$ and $t\mapsto n^{n}/(n-1)^{n-1}$ with homogenised irreducible polynomials $T,S$ and $(n-1)^{n-1}T-n^{n}S$. As explained on [24, p.283], the residue fields at $t\mapsto 0$ and tmaps-$t\mapsto\infty$ are $\mathbb{Q}(\mu_{n})$ and $\mathbb{Q}(\mu_{n-1})$, and the residue field at $t\mapsto n^{n}/(n-1)^{n-1}$ is the splitting field of the polynomial
@@
Proposition A.2, with a choice of $R,l,k$ to be specified later, produces infinitely many coprime pairs $(s_{0},t_{0})\in\mathbb{Z}^{2}$ satisfying (A.2)–(A.4). According to [24, Theorem 3.1 and Theorem 3.2], for large enough $|s_{0}|,|t_{0}|$, the only primes ramified in the specialisation $E_{\frac{t_{0}}{s_{0}}}/\mathbb{Q}$ are the primes $|s_{0}|,|t_{0}|$ and $|(n-1)^{n-1}t_{0}-n^{n}s_{0}|$, and moreover the decomposition groups at these primes are cyclic, and more precisely generated (in this order) by an $(n-1)$-cycle, an $n$-cycle and a transposition respectively.
This gives infinitely many extensions $E_{\frac{t_{0}}{s_{0}}}/\mathbb{Q}$ that satisfy conditions (2) and (3) of Theorem A.1. The proof of [24, Theorem 5.5] provides instructions on how to choose $R,l,k$ in order to ensure that condition (3) holds as well, if only the primes in $M_{2}$ are sufficiently large.
Moreover, as explained at the start of the proof of [24, Theorem 5.5], we get condition (1) for free by enlarging the set $M_{2}$ in order to ensure that all conjugacy classes of $\operatorname{S}_{n}$ occur amongst the $C_{p}$.
@@
It is furthermore natural to use extensions such as the ones obtained in the proof of Theorem A.1 to try to solve certain embedding problems and in particular to construct locally cyclic Galois extensions whose Galois group $G$ is a central extension of $\operatorname{S}_{n}$. This was done in [24, Theorem 5.6] for the case $n=5$ and for any $G$ which is a central extension of $\operatorname{S}_{5}$ in which the transpositions lift to elements of order $2$. The following observation shows that this approach does at least not generalize to arbitrary $n$ using the $\operatorname{S}_{n}$ extensions obtained here. Denote by $2.\operatorname{S}_{n}^{+}$ the one of the two stem covers of $\operatorname{S}_{n}$ in which the transpositions lift to elements of order $2$. It is then well-known (see, e.g., [19]) that the involutions of $\operatorname{S}_{n}$ which lift to involutions in $2.\operatorname{S}_{n}^{+}$ are exactly the products of $4j$ or $34j-$ 3 disjoint transpositions ($j\in\mathbb{N}$). On the other hand, the extensions constructed in Theorem A.1 arise via trinomials $X^{n}-\frac{t_{0}}{s_{0}}(X-1)$, and those have at most three real roots, meaning that complex conjugation acts as an involution with at most three fixed points. Whenever $n\equiv 6,7\pmod{8}$, such an involution hence lifts to elements of order $4$ in $2.\operatorname{S}_{n}^{+}$, meaning that the $\operatorname{S}_{n}$ extensions thus constructed do not embed into $2.\operatorname{S}_{n}^{+}$-extensions due to an obstruction at the archimedean prime.

### v3

@@ 77 $\frac{914838624}{353220125}\mathscr{A}_{7}^{4}$
Apart from the application of the new Proposition A.2, this is essentially the same as the proof of [25, Theorem 5.5], so we will be brief and focus on what is different.
We consider the splitting field $E/\mathbb{Q}(t)$ of the polynomial $f(t,X)=X^{n}-t(X-1)$, which is ramified over the places $t\mapsto 0,t\mapsto\infty$ and $t\mapsto n^{n}/(n-1)^{n-1}$ with homogenised irreducible polynomials $T,S$ and $(n-1)^{n-1}T-n^{n}S$. As explained on [25, p.283], the residue fields at $t\mapsto 0$ and $t\mapsto\infty$ are $\mathbb{Q}(\mu_{n})$ and $\mathbb{Q}(\mu_{n-1})$, and the residue field at $t\mapsto n^{n}/(n-1)^{n-1}$ is the splitting field of the polynomial
@@
Proposition A.2, with a choice of $R,l,k$ to be specified later, produces infinitely many coprime pairs $(s_{0},t_{0})\in\mathbb{Z}^{2}$ satisfying (A.2)–(A.4). According to [25, Theorem 3.1 and Theorem 3.2], for large enough $|s_{0}|,|t_{0}|$, the only primes ramified in the specialisation $E_{\frac{t_{0}}{s_{0}}}/\mathbb{Q}$ are the primes $|s_{0}|,|t_{0}|$ and $|(n-1)^{n-1}t_{0}-n^{n}s_{0}|$, and moreover the decomposition groups at these primes are cyclic, and more precisely generated (in this order) by an (n1)(n-1)-cycle, an $n$-cycle and a transposition respectively.
This gives infinitely many extensions $E_{\frac{t_{0}}{s_{0}}}/\mathbb{Q}$ that satisfy conditions (2) and (3) of Theorem A.1. The proof of [25, Theorem 5.5] provides instructions on how to choose $R,l,k$ in order to ensure that condition (4) holds as well, if only the primes in $M_{2}$ are sufficiently large.
Moreover, as explained at the start of the proof of [25, Theorem 5.5], we get condition (1) for free by enlarging the set $M_{2}$ in order to ensure that all conjugacy classes of $\operatorname{S}_{n}$ occur amongst the $C_{p}$.
@@
It is furthermore natural to use extensions such as the ones obtained in the proof of Theorem A.1 to try to solve certain embedding problems and in particular to construct locally cyclic Galois extensions whose Galois group $G$ is a central extension of $\operatorname{S}_{n}$. This was done in [25, Theorem 5.6] for the case $n=5$ and for any $G$ which is a central extension of $\operatorname{S}_{5}$ in which the transpositions lift to elements of order $2$. The following observation shows that this approach does at least not generalize to arbitrary $n$ using the $\operatorname{S}_{n}$ extensions obtained here. Denote by $2.\operatorname{S}_{n}^{+}$ the one of the two stem covers of $\operatorname{S}_{n}$ in which the transpositions lift to elements of order $2$. It is then well-known (see, e.g., [19]) that the involutions of $\operatorname{S}_{n}$ which lift to involutions in $2.\operatorname{S}_{n}^{+}$ are exactly the products of $4j$ or 4j34j-3 disjoint transpositions ($j\in\mathbb{N}$). On the other hand, the extensions constructed in Theorem A.1 arise via trinomials $X^{n}-\frac{t_{0}}{s_{0}}(X-1)$, and those have at most three real roots, meaning that complex conjugation acts as an involution with at most three fixed points. Whenever $n\equiv 6,7\pmod{8}$, such an involution hence lifts to elements of order $4$ in $2.\operatorname{S}_{n}^{+}$, meaning that the $\operatorname{S}_{n}$ extensions thus constructed do not embed into $2.\operatorname{S}_{n}^{+}$-extensions due to an obstruction at the archimedean prime.